A Stone has an initial velocity of 200 ft per see up to the right at a slope of 4 to 3. The components of acceleration are constant at a_x=-12 ft per second² and a_y=-20 ft per second². Compute the radius of curvature at the start and at the top of the path


1
Expert's answer
2021-03-18T03:37:45-0400

V0 = 200 fps ax = -12 fps2; ay = -20 fps2

Since we are given that the acceleration is constant, we can use

vx = v0x + axt

vy = v0y + ay

tan θ = ay/ax = -20/(-12) = 5/3

θ = tan-1(5/3) = 59 degrees

v0x = v0*cos θ = 200 fps * cos(59) = 103 fps

v0y = v0*sin θ = 200 fps * sin(59) = 171 fps

vx = 103 - 12t

vy = 171 - 20t